OpenStudy (anonymous):

Find the solution of y = 6x + 1 for x = 5

OpenStudy (anonymous):

Plug 5 for x in the equation y=6*5+1=30+1=31

OpenStudy (campbell_st):

substitute x with 5 \[y = 6\times5 + 1\]